# CLIP Won’t Learn Object-Attribute Binding from Natural Data and Here is Why

### v1

@@ 4 Results
Table 1: Attribute-Object binding in optimal setup. Here, Realistic Data refers to setup with properties determined from real data. Ideal Data refers to the best performing dataset properties as found in section 4.2. We ablate the Realistic and Ideal data setting with small and Ideal (large) batch size and model size. As seen below, the ideal data is much more important for high binding. This shows that the data properties limit binding, not model size or batch size.
Setup Attribute Binding Accuracy
@@
Ideal 256 256 99.99 98.55 98.01 99.93 98.56 99.46
4.1 CLIP Only Learns Attribute-Object Binding with Ideal Data
Can CLIP even learn object-attribute binding? Does the architecture and loss only promote learning a bag-of-words (BOW) representation or actively encourage it? As detailed above, a large batch size encourages object-attribute binding as more hard negatives are in the batch. We can explore this effect in MADMAN with relatively small batch sizes since the number of factors of variation is low.
As can be seen in Tab. 1, neither training with very large batch size, nor training a large model, nor their combination enables binding. We conclude that under realistic data settings, CLIP models are unlikely to learn binding. So either CLIP models can not learn binding, or properties of the data inhibit them. To answer this, we search for an ideal training data setup. As detailed in section 4.2, we find that changing the data setup has a significant effect on binding accuracy, leading to almost perfect binding (see Tab. 1). We conclude that CLIP models are able to learn object-attribute binding but properties of the data inhibit this.
4.2 What Limits the Object Binding in CLIP Models?
The previous section shows that CLIP models can learn object-attribute binding under optimal data conditions. But which properties of the data are important? How do they individually influence the binding? Can we find factors that we could change on real data? To answer these questions we investigate the relationship of 4 data properties on the binding performance below. Unless otherwise mentioned, we set Multi-object image probability to $1.0$, Multi-object caption probability to $1.0$, the expected number of Attributes per object in caption to $1.8$ and Saliency bias to 0.
What’s the influence of the ratio of the ratio of multi-object images on binding?
(a) Multi-object image probability
(b) Multi-object caption probability
Figure 3: Influence of (a) Multi-object image probability and (b) Multi-object caption probability on binding accuracy. Having more images and captions containing multiple objects improves binding.
To answer this we vary Multi-object image probability. As can be seen in Fig. 3(a), object-attribute binding increases steadily for all attributes.
What’s the influence of the ratio of multi-object captions on binding? Similarly, the number of objects mentioned in the caption likely influences binding. Fig. 3(b) reveals that small ratios are already sufficient to learn binding for a subset of the attributes.
What’s the influence of the attributes per object in caption on binding?
(a) Attributes per object in caption
@@ 4.2 What Limits the Object Binding in CLIP Models?
Figure 4: Influence of (a) Attributes per object in caption and (b) Saliency bias on binding accuracy. (a) Too few and too many Attributes per object in caption hamper the binding accuracy. For large Attributes per object in caption binding is not necessary anymore, as BOW is completely sufficient to discriminate images due to combinatorial explosion. (b) Saliency bias hinders binding, especially at higher values, which is the case for natural data. This is likely because it provides a shortcut for the model to minimize contrastive loss by simply representing the salient object(s) without binding them.
As visible in Fig. 4(a), the Attributes per object in caption have a strong influence on binding. Interestingly, for expected number of attributes larger than $3.5$ the binding accuracy starts to decrease again. We suspect that this is due to the larger discriminativeness of a BOW representation with many concepts due to the combinatorial explosion. When reaching a certain threshold value the network stops learning binding as it is not necessary to learn it. With low expected value the probability of both objects having an attribute in the caption is low and, therefore, the learning signal to learn binding weak.
@@
What’s the influence of the saliency bias of captioners on binding? Humans typically agree on which object(s) are the most important (or salient) in an image [3, 4]. Saliency might be problematic for binding as it provides shortcut-opportunities to minimize the contrastive loss, for instance by representing primarily the attributes of the salient object without binding them  [^1] Note that the text encoder has also information which object is likely the salient one, as the salient object is always mentioned first in the caption..
@@
Indeed we find that small levels of saliency already have a negative effect, while levels above 0.75 are detrimental (Fig. 4(b)). As humans tend to take pictures of objects and not random images and tend to describe the salient objects with higher likelihood, we expect high saliency bias values on real data. Indeed, on the manually labeled images of CC12M we identified for more than 90% of the images at least one salient object (see Appendix, section 5.2), suggesting that saliency is likely one of the main limiting factors for CLIP models to learn binding.
@@
Color often behaves differently. Interestingly, we observe that color is bound only for high Multi-object image probability (Fig. 3(a)) and Multi-object caption probability (Fig. 3(b)), despite being easily recognizable (see Appendix, section 5.4) and having most discriminative power among the attributes, e.g., knowing that there is a green 6 in the image distinguishes it from more images than knowing there is a small 6, as there are more colors than scales. So why does it bind so poorly? We suspect that this high discriminative power is exactly the reason: Having access to color(s) and object class(es) alone is often sufficient to discriminate from all other samples in the batch, making it unnecessary to bind color.
Also for the number of attributes per object in the caption (Fig. 4(a)) the binding of color peaks at much higher levels than for attributes with fewer possible values. But why does it learn binding for color later? As the Attributes per object in caption increases, the probability of color appearing in the caption increases. Knowing the color(s) is often already sufficient to discriminate from all other samples in a batch. As color is also very easy to recognize, the model does not learn high recognition of other attributes (see Appendix, section 5.4) and relies more on color. This over-reliance on color seems to encourages the model to learn binding of colors instead of recognizing other attributes. These findings highlight that there is a complex interplay between attribute frequency, difficulty to recognize an attribute and discriminability.
[…]
@@
Figure 5: Experimental results are independent of most important hyper-parameters of the experimental setup. We test various combinations of the embedding size and batch size. We observe the same trends for all data properties as found in our main setting, showing that our results are not an artifact of a specific experimental setup.
@@ 4.3 Validating our Experimental Setup
To validate that our findings are not just an artifact of poor choices for the batch size and model size, we repeat the main experiments and vary the embedding-size of the models and the batch size. By changing the embedding-size we test whether the model size limited binding. By changing the batch size, we test whether the hard-negative ratio (which is unknown for real data) was underestimated. Fig. 5(b) reveals that our findings are independent of the batch size. This also holds for a bigger model as seen in Fig. 5(c). Changing the batch size leads to slightly different results, but the trends are very similar. A similar observation can be made for Fig. 5(a). We observe small changes in performance, however, the overall trends remain the same. Combined, these results validate our experimental setting and findings as not being an artifact of the base setup.
@@ 4.4 We find the Same Relationships Out-of-distribution
(a) Multi-object image probability
(b) Multi-object caption probability
(c) Attributes per object in caption
@@
Figure 6: Impact of the data properties on the binding for out-of-distribution combinations. It follows similar trends as in-distribution (denoted by the dashed line).
CLIP models have been shown to learn object-attribute compositional representations [24]. However, in a synthetic setting with few factors of variation a model can just treat each object and attribute combination as an individual class, instead of learning such a compositional representation. While the compositional representation can generalize to unseen class & attribute combinations, the non-compositional representation can not. If our models learn non-compositional representations we can not expect our results to transfer to large scale CLIP training. To test this, we test our models on our OOD test set with unseen class & attribute combinations. We find the same trends also in OOD (see Fig. 6), which confirms that our model learns a similar compositional representation as CLIP models trained on real data and, thus, confirms the validity of our study.
@@ 4.5 Are Hard Negatives Enough to Solve the Binding Problem?
CLIP 49.95 54.23 52.04 51.00 53.87 50.31
CLIP (+Ideal data) 94.70 91.27 61.99 92.47 66.40 90.12
@@
Finally, we explore how well NegCLIP performs on the variant of MADMAN following realistic data properties. Does NegCLIP alone resolve the binding problem? As Tab. 2 reveals, NegCLIP indeed improves binding over CLIP, but does not get close to CLIP trained in the ideal data setup. Thus, NegCLIP alone is unlikely to resolve the issue and better data is necessary.
@@ 5 Conclusion
In this work, we showed that CLIP struggles to learn object-attribute binding even when training under favorable conditions. Via a rigorous analysis using our synthetically generated dataset we showed that specific data properties common in web-scraped image-caption datasets are limiting CLIP’s object-attribute binding. In particular, we find that having too few images showing, and too few captions describing multiple objects limits binding. More importantly, we show that too few, as well as too many attributes mentioned in the caption per object hurt binding. Most interestingly, we find that the saliency bias plays a critical role and, at levels that are common on real data has detrimental effects on the binding performance. Indeed, if the saliency bias is too high the influence of the other properties diminishes. We verify our experiments thoroughly, by changing model parameters and testing in an OOD setup.
Limitations. This work shows the relationship of the aforementioned data properties to binding accuracy of CLIP models, however, our studies are limited to a synthetic setup. We expect to find the same relationship on real data, however, obtaining the required information to perform these experiments on real data is prohibitively expensive.
Future work should use the insights provided by this work to guide filtering or re-captioning approaches. A promising re-captioning approaches is to occasionally drop the salient object from the caption. Another promising direction is to explore the optimal number of attributes per object in the caption on real data.
@@ Acknowledgments
We would like to thank Simon Schrodi for helpful comments on the draft. Funded by the German Federal Ministry for Economic Affairs and Energy within the project “NXT GEN AI METHODS” (19A23014R), by the German Research Foundation (DFG) - 417962828, 539134284 and – SFB 1597 – 499552394.
@@ 5.1 Dataset Creation and Model Setup
We introduce MADMAN, a synthetic dataset designed to systematically investigate how dataset properties affect attribute binding. The dataset enables precise control over three key aspects: the number of objects per image, the completeness of object descriptions in captions, and the density of attribute descriptions per object.
@@
Table 3: Transformations applied on the digits. The categories act as the attributes and the type within each category that we apply to the digit is the value of the attribute. We thus have $(3*2*2*2*3*7)=504$ possible combinations of attribute values.
@@
To create a multi-object setup, we create a 3x3 grid, consisting of 9 cells of 28x28 pixels each, and fill two of them with samples from MAD above to create a multi-object image of size 96x96.
@@
[…]
@@
We generate captions by treating attributes and digit classes as individual tokens. For each object, we create a noun phrase with the attributes appearing before the digit class. Next, we combine the noun phrases using and as the conjunction token. All tokens in the caption are separated by the space character. Note: as listed in Tab. 3, we have 19 unique attribute values for the six attributes. With the ten object classes and four special tokens (and, <pad>, <start>, <end>), our vocabulary size is 33 in total. The generation process is as follows:
@@
[…]
@@
    • We vary the number of attributes ($n_{a}$) used to describe an object by sampling it from a categorical distribution (see Fig. 7).
@@
$o_{i}=\begin{cases}2&\text{with probability }p(\text{multi-obj-img}),\\ 1&\text{with probability }1-p(\text{multi-obj-img}).\end{cases}$
How often are the objects described in the caption?
[…]
$o_{T}=\begin{cases}2&\text{with probability }$p(\text{multi-obj-cap})$,\\ 1&\text{with probability }1-$p(\text{multi-obj-cap})$\end{cases}$
@@
One data property we are interested in is the number of attributes associated with an object that are included in the captions. As seen in Fig. 7, we define a categorical distribution for this number $n_{a}$ over the possible number of attributes that can be included i.e. [0, 6]. The different distributions simulate having less / more of the attributes being included per object description in the caption.
@@
Figure 9: Samples from the MADMAN dataset. The images are generated from transformed MNIST instances while the captions is created describes the attributes and the object class. There can be variable number of objects in the image, objects in the caption, and attributes described per object.
@@
Table 4: Comparison of architecture parameters between our model TinyCLIP, used as the base model for the experiments, and openai/clip-vit-base-patch32. Due to the simple nature of the dataset we are working with (e.g. vocab_size is three orders of magnitude smaller), we work with a smaller model. But as seen in section 4.3 the results also hold for a bigger model.

### v2

@@ 4 Results
Table 1: Object-Attribute binding in realistic and ideal setup. Here, Realistic Data refers to a setup with properties determined from real data. Ideal Data refers to the best performing dataset properties as found in Section˜4.2. We ablate the Realistic and Ideal data setting with small and Ideal (large) batch size and model size. It can be seen that ideal data is more important for high binding. Thus, data properties limit binding, not model size or batch size.
Setup Binding Accuracy
@@
Ideal 256 256 100.0 98.48 98.18 99.93 98.64 99.43
4.1 CLIP Only Learns Object-Attribute Binding with Ideal Data
Can CLIP even learn object-attribute binding? Does the architecture and loss prevent CLIP from learning a non Bag-of-words (BOW) representation? Is maybe a large enough batch size sufficient to encourage CLIP to learn binding via hard negatives? We can explore these questions using MADMAN with relatively small batch sizes since the number of factors of variation is low.
As can be seen in the top part of Tab.˜1, neither training with very large batch size, nor training a large model, nor their combination enables binding. We conclude that CLIP models are unlikely to learn binding when training on data that exhibits realistic values for the four data properties (defined in Section˜3.1.1). So either CLIP models can not learn binding, or properties of the data inhibit it. To find out which of these two explanations is true, we optimize the training data setup by changing the four data properties. As detailed in Section˜4.2, we find that changing the data setup has a significant influence on binding accuracy, leading to almost perfect binding (see Tab.˜1). We conclude that CLIP models are able to learn object-attribute binding but properties of the data inhibit this.
4.2 What Limits the Object-Attribute Binding in CLIP Models?
The previous section shows that CLIP models can learn object-attribute binding under optimal data conditions. But which properties of the data are important? How do they individually influence the binding? Can we find factors that we could change on real data? To answer these questions, we individually investigate the influence of the four data properties defined in Section˜3.1.1 on the binding performance. Unless otherwise mentioned, we set Two-object-in-image-probability to $1.0$, Two-object-in-caption-probability to $1.0$, the expected number of Attributes-per-object-in-caption to $1.8$ and Saliency bias to 0. For all experiments below, we vary one data property and keep the others constant. By varying one property we create different variants of MADMAN. For each of these versions we train 3 CLIP models from scratch and report the average binding accuracy.
4.2.1 What’s the influence of the ratio of multi-object images on binding?
(a) Two-object-in-image-probability
(b) Two-object-in-caption-probability
Figure 3: Influence of (a) Two-object-in-image-probability and (b) Two-object-in-caption-probability on binding accuracy on MADMAN. Having more images and captions containing multiple objects improves binding.
To answer this we vary Two-object-in-image-probability. As can be seen in Fig.˜3(a), object-attribute binding increases steadily for all attributes.
4.2.2 What’s the influence of the ratio of two-object captions on binding?
Similarly, the number of objects mentioned in the caption likely influences binding. Fig.˜3(b) reveals that a small ratio of captions containing two objects are already sufficient to learn binding for a few attributes. Increasing the number of captions containing two objects leads to consistent improvements of binding.
4.2.3 How does the number of attributes per object in the caption influence binding?
(a) Attributes-per-object-in-caption
@@ 4.2.3 How does the number of attributes per object in the caption influence binding?
Figure 4: Influence of (a) Attributes-per-object-in-caption and (b) Saliency bias on binding accuracy on MADMAN. (a) Too few and too many Attributes-per-object-in-caption hamper the binding accuracy. For large Attributes-per-object-in-caption binding is not necessary anymore, as BOW is sufficient to discriminate images due to combinatorial explosion. (b) Saliency bias hinders binding, especially at higher values, which is the case for natural data. This is likely because it provides a shortcut for the model to minimize contrastive loss by simply representing the salient object(s) without binding them.
As can be seen in Fig.˜4(a), the Attributes-per-object-in-caption have a strong influence on binding. Interestingly, for Attributes-per-object-in-caption larger than $3.5$ the binding accuracy starts to decrease again. We suspect that this is due to the larger discriminativeness of a BOW representation with many concepts due to the combinatorial explosion: If enough attributes are present, a BOW representation is sufficient to discriminate between all samples in a batch (compare Fig.˜1(b)) and thus the model does not need to learn binding. In the other extreme, low Attributes-per-object-in-caption prevents binding from being learned, as cases with two attributes where binding is necessary, are too rare.
4.2.4 What’s the influence of the saliency bias of captioners on binding?
@@
Humans typically agree on which object(s) are the most important (or salient) in an image [4, 5]. Saliency might hamper binding as it provides shortcut-opportunities to minimize the loss, e.g., by representing primarily the attributes of the salient object without binding them. [^1] Note that the text encoder also has information on which object is likely the salient one, as the salient object is always mentioned first in the caption. Indeed we find that already small levels of saliency have a negative effect, while levels above 0.75 are detrimental (Fig.˜4(b)).
@@
But can we really expect more than 75% of image-caption pairs to express saliency bias? We expect this to be the case as humans tend to take pictures of objects instead of random images and tend to describe the salient objects with higher probability. To verify this, we use our manually labeled images of CC12M. Indeed, we identified at least one salient object for more than 90% of the images (see Appendix, Section˜5.2). In our experiments we observe very low binding for these saliency levels (Fig.˜4(b)), suggesting that saliency is likely one of the main limiting factors for CLIP models to learn binding. We even find that the influence of other factors diminishes when saliency is at a natural level (see Tab.˜6). Thus saliency might be a key limiting factor also on real data.
4.2.5 Color often behaves differently.
@@
Interestingly, we observe that the attribute color binds poorly for most values of Two-object-in-image-probability (Fig.˜3(a)) and Two-object-in-caption-probability (Fig.˜3(b)), despite being easy to recognize (see Appendix, Section˜5.5). But why does it bind so poorly? Note that the color attribute can take 7 different values, while all other attributes have fewer possible values. As a result, color has most discriminative power among the attributes, e.g., knowing that there is a green 6 in the image distinguishes it from more images than knowing there is a small 6. Thus, if color(s) and object(s) are in the caption, representing these two factors is often sufficient to distinguish this image/caption from all other captions/images in a batch, making binding unnecessary. Only when colors of more than one object are frequently in the caption, learning the color-object binding reduces the loss (relatively) enough s.t. the model eventually learns binding.
The ideal data setting. Finally, we use the previous experiments to select the best data setup for each of the properties. We end up with: no saliency bias, $p(\text{multi-obj-img})=1$, $p(\text{multi-obj-cap})=1$ and expected number of Attributes-per-object-in-caption = 3.5. As shown in Tab.˜1, combining the optimal setups for all our data properties indeed leads to high binding and combining the best setups for each parameter leads to best binding performance (see Appendix, Section˜5.3). Tab.˜1 reveals that ideal data alone leads to high object-attribute binding for 4 out of 6 attributes. Using ideal data, a large model and a large batch size results in almost perfect binding for all attributes.
@@
Figure 5: Experimental results are independent of most important hyper-parameters of the experimental setup on MADMAN. We test various combinations of the embedding size and batch size. We observe the same trends for all data properties as found in our main setting, showing that our results are not an artifact of a specific experimental setup.
@@ 4.3 Validating our Experimental Setup
The results presented above could be an artifact of an unlucky choice of the batch size or model size. To rule out that our findings are just an artifact of poor choices for the batch size and model size, we repeat the experiments and vary the batch size and the embedding size of the models. As detailed above, the batch size directly controls the number of hard negatives in the loss, thus testing different batch sizes can be interpreted as changing the hard negative ratio in a batch. By changing the embedding size we test whether the model size limited binding in our experiments. Fig.˜5(a) shows that our results are not an artifact of the embedding size. We observe small differences, however, the trends remain the same. Similarly, Figs.˜5(b) and 5(c) reveal that our findings are independent of the batch size. Changing the batch size leads to small differences, but the trends remain consistent. Combined, these results validate our experimental setting and findings as not being an artifact of the base setup. The only exception of this is the Two-object-in-caption-probability, where the effect partially diminishes with increasing embedding size and increasing batch size (Figs.˜5(a), 5(b) and 5(c)).
@@ 4.4 We find the Same Relationships Out-of-distribution
(a) Two-object-in-image-probability
(b) Two-object-in-caption-probability
(c) Attributes-per-object-in-caption
@@
Figure 6: Impact of the data properties on the binding for out-of-distribution combinations on MADMAN-OOD. It follows similar trends as in-distribution (denoted by the dashed line).
CLIP models have been shown to learn object-attribute compositional representations [27], meaning that the concept “yellow submarine” is encoded roughly by the sum of the “yellow” vector and the “submarine” vector. However, in a synthetic setting with few factors of variation a model can just treat each object- attribute combination as an individual class, instead of learning such a compositional representation. If our models learn non-compositional representations we can not expect our results to transfer to large scale CLIP training. To test this, we test our models on our OOD test set with unseen object-attribute combinations. A compositional representation can generalize to unseen object-attribute combinations, while a non-compositional representation can not. We also find the same trends for OOD (see Fig.˜6). Thus, our model learns a similar compositional representation as CLIP models trained on real data, which confirms the validity of our study.
@@ 4.5 Are Hard Negatives Enough to Solve the Binding Problem?
CLIP 50.47 54.47 52.34 50.86 53.26 50.05
CLIP (+Ideal data) 94.66 91.28 61.93 92.44 66.43 90.02
@@
Previous works tried to resolve binding by changing the training process, for instance by training on hard negatives [29]. The previous sections show that data properties of the training data have a large influence on the binding performance of CLIP models. Naturally, the question arises whether a hard negative training regime, as proposed for NegCLIP [29], can be used to resolve the problems induced by the data properties. To this end, we explore how well NegCLIP performs on the variant of MADMAN following realistic data properties. As Tab.˜2 reveals, NegCLIP indeed improves binding over CLIP, but does not get close to CLIP trained in the ideal data setup. Thus, NegCLIP alone is not enough to resolve the issues caused by the data properties and thus is unlikely to resolve the issue on real data. Our results indicate that better training data might be necessary to resolve the binding problem of CLIP.
@@ 5 Conclusion
In this work, we showed that CLIP struggles to learn object-attribute binding even when training under favorable conditions. Via a rigorous analysis using our synthetically generated dataset we showed that specific data properties common in web-scraped image-caption datasets are limiting CLIP’s object-attribute binding. In particular, we found that having too few images showing, and too few captions describing multiple objects limits binding. More importantly, we showed that too few and too many attributes mentioned in the caption per object hurt binding. Most interestingly, we found that the saliency bias plays a critical role and, at levels common on real data, has a detrimental effect on the binding performance. Furthermore, we found that when the saliency bias is too high, the influence of the other properties diminishes. We verified our experiments thoroughly, by changing model parameters and testing in an OOD setup, where we found the same trends. In summary, our controlled experiments on synthetic data provide evidence that the data properties common for natural image-caption datasets have a negative influence on CLIP’s binding performance. To obtain high binding performance of CLIP models on natural data it might be necessary to change some of the data properties of the training data.
Limitations. This work shows the relationship of the aforementioned data properties to binding accuracy of CLIP models; however, our studies are limited to a synthetic setup. We expect to find the same relationship on real data but obtaining the required information to perform these experiments on real data is prohibitively expensive.
Future work should use the insights provided by this work to guide filtering or re-captioning approaches. A promising re-captioning approach is to occasionally drop the salient object from the caption. Another promising direction is to explore the optimal number of attributes per object in the caption on real data.
@@ Acknowledgments
We would like to thank Simon Schrodi for helpful comments on the draft. Funded by the German Federal Ministry for Economic Affairs and Energy within the project “NXT GEN AI METHODS" (19A23014R), by the German Research Foundation (DFG) - 417962828, 539134284 and – SFB 1597 – 499552394.
@@ 5.1 Dataset Creation and Model Setup
We introduce MADMAN, a synthetic dataset designed to systematically investigate how dataset properties affect object-attribute binding. The dataset enables precise control over four key aspects: the number of objects per image, the completeness of object descriptions in captions, the density of attribute descriptions per object, and saliency bias.
@@
Table 3: Transformations applied on the digits. The transformation categories act as the attributes and the type within each category that we apply to the digit is the value of the attribute. We thus have (322237)=504(3*2*2*2*3*7)=504 possible combinations of attribute values.
@@
To create a multi-object setup, we create a 3x3 grid, consisting of 9 cells of 28x28 pixels each, and fill two of them with samples from MAD above to create a two-object image of size 96x96.
@@
    Select $o_{i}$=(1 or 2) objects to add to the image. We vary this using a parameter $p(\text{two-obj-img})$ that determines whether the second object is included or not.
@@
We generate captions by treating attributes and digit classes as individual tokens. For each object, we create a noun phrase with the attributes appearing before the digit class. Next, we combine the noun phrases using and as the conjunction token. All tokens in the caption are separated by the space character. Note: as listed in Tab.˜3, we have 19 unique attribute values for the six attributes. With the ten object classes and four special tokens (and, <pad>, <start>, <end>), our vocabulary size is 33 in total. The generation process is as follows:
@@
    Select $o_{t}$=(1 or 2) objects to add to the caption. We vary this by a parameter $p(\text{two-obj-cap}|\text{two-obj-img}=\text{True})$ that determines whether the second object is included or not. Note: if there is only one object in the image, we can take only that one object in the caption.
@@
    • We vary the number of attributes ($n_{a}$) used to describe an object by sampling it from a categorical distribution (see Fig.˜7).
@@
Real-world images can contain $o_{i}\geq 1$ objects. We capture this in our minimal setting by having $o_{i}\in\{1,2\}$ controlled by Two-object-in-image-probability $p(\text{two-obj-img})$.
$o_{i}=\begin{cases}2&\text{with probability }p(\text{two-obj-img}),\\ 1&\text{with probability }1-p(\text{two-obj-img}).\end{cases}$
How often are the objects described in the caption?
Given an image with multiple objects (in our case, two), a caption may or may not describe all of them. We capture this aspect in our setting by having $o_{T}\in\{1,2\}$ controlled by Two-object-in-caption-probability $p(\text{two-obj-cap})$.
@@
One data property we are interested in is the number of attributes associated with an object that are included in the captions. As seen in Fig.˜7, we define a categorical distribution for this number $n_{a}$ over the possible number of attributes that can be included i.e. [0, 6]. The different distributions simulate having less / more of the attributes being included per object description in the caption.
@@
Figure 9: Samples from the MADMAN dataset. The images are generated from transformed MNIST instances while the captions are created by chaining object (digit) name and the set of attributes (augmentations applied). There can be variable number of objects in the image, objects in the caption, and attributes described per object.
@@
Table 4: Comparison of architecture parameters between our model TinyCLIP, used as the base model for the experiments, and openai/clip-vit-base-patch32. Due to the simple nature of the dataset we are working with (e.g. vocab_size is three orders of magnitude smaller), we work with a smaller model. But as seen in Section˜4.3 the results also hold for a bigger model.